# MFIT : Multi-FIdelity Thermal Modeling for 2.5D and 3D Multi-Chiplet Architectures

### v2

@@ Abstract
Rapidly evolving artificial intelligence and machine learning applications require ever-increasing computational capabilities, while monolithic 2D design technologies approach their limits. Heterogeneous integration of smaller chiplets using a 2.5D silicon interposer and 3D packaging has emerged as a promising paradigm to address this limit and meet performance demands. These approaches offer a significant cost reduction and higher manufacturing yield than monolithic 2D integrated circuits. However, the compact arrangement and high compute density exacerbate the thermal management challenges, potentially compromising performance. Addressing these thermal modeling challenges is critical, especially as system sizes grow and different design stages require varying levels of accuracy and speed. Since no single thermal modeling technique meets all these needs, this paper introduces MFIT, a range of multi-fidelity thermal models that effectively balance accuracy and speed. These multi-fidelity models can enable efficient design space exploration and runtime thermal management. Our extensive testing on systems with 16, 36, and 64 2.5D integrated chiplets and 16$\times$3 3D integrated chiplets demonstrates that these models can reduce execution times from days to mere seconds and milliseconds with negligible loss in accuracy.
I Introduction
@@
Massive data from different modalities, including text, images, video, and speech, are continuously produced by various sensors. At the same time, increasingly complex artificial intelligence (AI) and machine learning (ML) algorithms process this data to enable new applications that were previously impractical. This trend dictates the design of large-scale chips with high memory and compute capabilities, offering a high degree of parallelism [1, 2]. Traditional 2D chip design and packaging technologies cannot sustain this need due to the low yield of large monolithic planar chips and the corresponding increase in fabrication cost [3]. Therefore, new design approaches are required to meet the increasing demand for computing power and memory capacity [1].
@@
2.5D and 3D chiplet-based architectures have emerged as promising alternatives to traditional monolithic 2D chips due to their lower fabrication costs [4, 5, 6]. Compared to conventional monolithic systems, chiplet-based systems integrate multiple small pre-fabricated chips (chiplets) on a silicon interposer, which facilitates data exchange, as illustrated in Figure 1(a). 3D packaged systems expand on this approach by stacking multiple chiplets vertically and connecting them with vertical vias, creating a more compact system as illustrated in Figure 1(b). The smaller size of these chiplets enables a higher yield and lower overall manufacturing cost than traditional monolithic dies [7]. Additionally, this modular approach facilitates scaling the system sizes and enables heterogeneous integration of different chiplet types, e.g., memory, processing, and processing-in-memory chiplets. Hence, emerging 2.5D and 3D architectures enable a new cost-effective avenue for compact scale-out implementations of various emerging compute- and data-intensive applications, including AI/ML. Indeed, these advantages have led to industrial adoption by companies including Intel [8, 9], AMD [10, 11, 12], and NVIDIA [4].
@@
(a) A 16 - 2.5D integrated chiplet based system. The magnified view on the right shows a detailed structure of a single chiplet.
@@
(b) A 16$\times$3 - 3D integrated chiplet based system. The magnified view on the right shows a detailed structure of three chiplets stack.
@@
Figure 1: 2.5D and 3D integrated chiplet systems considered in this work, showing the chiplets, interposer, and a portion of the substrate.
Thermal bottlenecks have long been a significant barrier to increasing the performance of computing systems. 2.5D and 3D integrated systems exacerbate this barrier due to their dense integration and unique physical structure [13]. In contrast to a monolithic chip, where heat is spread directly across the die, a 2.5D chiplet-based system conducts heat between different chiplets through the interposer and heat spreader. Likewise, heat also flows vertically between adjacent stacked chiplets in a 3D chiplet-based system.
@@
Figure 2: Summary of the multi-fidelity thermal models. (1) Fine-grained FEM models capture precise geometry but are too complex to simulate the entire chiplet-based system. (2) Abstracted FEM models are derived from the fine-grained model to simulate large-scale systems with negligible impact on accuracy. (3) Since abstract FEM models are still too slow for DSE, they are used to tune thermal RC circuit models. (4) Further abstraction reduces the execution time to milliseconds using DSS models created for specific system configurations, enabling runtime thermal management.
@@
The semiconductor chip design cycle spans multiple phases: system specification, architecture exploration, logic design, physical design and validation, fabrication, and post-silicon optimization/validation. Each phase has a unique set of design constraints and requirements. For example, lacking a test chip during the pre-silicon phases requires simulation and analytical models. Finite Element Method (FEM) simulations offer the most accurate approach for pre-silicon thermal analysis [14]. They can serve as a reference and enable heat flow studies to guide the design process. However, they are too slow for practical architecture and design space exploration (DSE). Modeling the package as a thermal RC (resistive-capacitive) network can significantly accelerate simulations with acceptable accuracy loss [15, 16]. Since each node in the thermal circuit corresponds to a specific location in the package, thermal RC models solve discretized versions of the FEM models in space. Hence, they enable thermally-aware DSE and optimization with a finite number of discrete hotspot nodes. However, the thermal resistance/capacitance values and the circuit topology must accurately reflect the chip geometry and material properties for reliable results. Since the thermal RC models solve continuous-time ordinary differential equations (ODEs), they have execution times in the order of seconds to minutes. Therefore, they cannot be used for runtime optimization tasks such as dynamic thermal and power management (DTPM). One can discretize them in the time domain with a given sampling period [17, 18]. The resulting discrete state-space (DSS) models significantly reduce runtime at the cost of further abstracting the model from the physical package. Consequently, they are applicable only to the specific configurations for which they are developed.
@@
There is a strong need for tools to accurately analyze the thermal behavior of 2.5D and 3D integrated systems and guide their design process. However, no single modeling technique can alone address the needs of all design phases. To address this much needed gap, this paper proposes MFIT, a multi-fidelity thermal modeling framework that synergistically exploits the strengths of each class of models (FEM, thermal RC, and DSS). We use this framework to produce a set of thermal models that can guide the entire design cycle, unlike a point solution that can serve a specific portion of the design process. The elements of this set not only cover complementary parts of the design cycle but support each other and produce consistent results. We first develop a fine-grained FEM model of the target package as a reference. Since it is slow and computationally expensive, we next judiciously design an abstracted version of this fine-grained FEM model to simulate an entire package in days while maintaining accuracy. To enable fast DSE, MFIT also incorporates thermal RC circuit models verified against the reference FEM models. Our thermal RC models run in the order of seconds while leading to less than 1.7C error, as summarized in Figure 2. Hence, they can be used for pre-silicon architectural optimization, such as mapping the workloads to chiplets, network-on-interposer design, and chiplet placement for 2.5D and 3D stacked systems. Finally, MFIT derives one more class of models by discretizing the thermal RC models, enabling runtime thermal management and large-scale DSE in the order of milliseconds. However, they work only for a specific sampling period and configuration. Hence, the parameters must be regenerated from the RC model if the target configuration changes. In summary, we obtain a set of multi-fidelity thermal models that guide and complement each other to cover all design phases.
@@
  • A family of open-source multi-fidelity thermal models that span a wide accuracy (reference to 1.7C) and speed (days to milliseconds) range,
@@
  • Open-sourced code for thermal RC and DSS models at github.com/AlishKanani/MFIT. Additionally, we plan to make our FEM models publicly accessible in the near future.
@@
The remainder of the paper is organized as follows. Section II and Section III discuss related work and background on FEM. Section IV presents the proposed multi-fidelity thermal modeling framework. Finally, Section V presents the experimental evaluation, and Section VI concludes the paper.
II Related Work
@@
2.5D and 3D integration-based systems are becoming mainstream due to higher performance and lower manufacturing costs than monolithic chips. Both domain-specific and general-purpose 2.5D and 3D architectures have been explored to date [11, 19, 20, 5, 4, 21, 12]. SIMBA is one of the first prototype multi-chip modules with 36 chiplets designed for inference with deep models [4]. Similarly, Floret is a data center-scale architecture for accelerating convolutional neural network (CNN) inferencing tasks by exploiting the dataflow knowledge [5]. Loi et al., [22] analyze the performance benefits of vertically integrated (3D) processor-memory hierarchy under thermal constraints. Similarly, Eckert et al. [21] consider processing-in-memory (PIM) architectures implemented using 3D die stacking. They study the thermal constraints across different processor organizations and cooling solutions to identify viable solutions. Our proposed open-source thermal models catalyze similar thermal analysis and optimization studies for 2.5D and 3D integrated systems.
@@
The most accurate and direct thermal evaluation approach is temperature measurements on a hardware system. It can be performed using thermal imaging [23, 24] or temperature sensors [25, 26]. However, the availability of the target hardware is a significant limitation. For example, large-scale 2.5D and 3D chiplet systems with tens of chiplets do not exist yet, while smaller prototypes and commercial systems provide limited insights applicable for larger systems [10, 27]. This limitation motivated FEM-based modeling as the most accurate way to analyze the heat flow and temperature [14]. Proprietary software, such as ANSYS Fluent [28] and COMSOL [29], are commonly used for FEM simulations. Since FEM suffers from computational cost, detailed FEM solutions are suitable only for small designs and validating analytical models [28]. For example, the authors of [30] employ FEM to simulate a two-chiplet system on an interposer. They employ abstracted FEM models for both $\mu$-bumps and C4 bumps to speed up the process, effectively reducing computational complexity before tackling the entire system.
@@
The computational overhead and impractically high execution time of FEM solvers motivate analytical models that enable rapid thermal evaluation in early design phases. The most common method involves constructing thermal RC networks and solving the corresponding system of ODEs. Popular thermal simulators such as HotSpot [15] leverage this method, focusing on the microarchitectural layout blocks to facilitate design space exploration and early-stage thermal-aware layout and placement. Similarly, 3D-ICE [16] models liquid cooling with microchannels embedded between silicon layers. PACT [31] also employs a similar methodology by utilizing SPICE tools as solvers, focusing on standard-cell-level thermal analysis for 2.5D systems. However, these tools are not fast enough for large-scale, thermally-aware DSE of multi-chiplet 2.5D and 3D integrated systems.
2.5D and 3D packages often involve materials with varying thermal conductivity across different directions. For example, the thermal conductivity of the C4 layer is higher in the vertical direction than in the lateral direction. The existing thermal models [15, 16, 31] do not account for these variations. Moreover, they assume a uniform grid size for all material layers (e.g., interposer, C4 bumps, chiplets). In contrast, the thermal RC models in our multi-fidelity set allow varying thermal conductivity across different directions and grid sizes for each layer and block.
@@
Architecture-level thermal RC models are well-suited for offline studies, such as architectural exploration, temperature sensor placement [32, 33], and thermal-aware chiplet placement [34]. However, DTPM requires a much faster temperature estimation time, in the order of milliseconds, for real-time temperature management [18, 35]. DSS models address this need by deriving a discrete-time linear time-invariant system that models the thermal dynamics at fixed locations as a function of the power consumption. For instance, TILTS [36] discretizes the power inputs to the chip over fixed time intervals to accelerate thermal simulations. Hence, it needs to be reproduced when the timing requirements or the underlying hardware configuration change. The speedup gain offsets the loss of the explicit connection to the hardware parameters (e.g., thermal conductance and capacitance) and generality.
@@
The results of FEM, thermal RC, or other thermal simulations can also be used to train physics-informed machine learning techniques to model heat transfer in integrated circuits to reduce the thermal modeling effort [37]. For example, a recent technique collects data from numeric simulations and trains a random forest model to predict the convection heat transfer coefficients for a nonlinear heat transfer problem [38]. Similarly, Hwang et al., [39] present closed-form models derived from numerical simulations for tapered micro-channels to analyze the heat transfer performance as a function of the channel geometry. In contrast to individual classes of thermal models, this work proposes a framework to produce a family of multi-fidelity thermal models for 2.5D and 3D chiplet-based systems. The specific set of models designed with this framework covers a wide range of accuracy and execution time trade-offs, making them suitable for different design phases. Additionally, they can be augmented by additional models, such as physics-informed ML models, with complementary accuracy and execution time trade-offs.
Figure 3: An illustration of the FEM pipeline.
III FEM for Thermal Analysis

### v4

@@ Abstract.
Rapidly evolving artificial intelligence and machine learning applications require ever-increasing computational capabilities, while monolithic 2D design technologies approach their limits. 2.5D/3D heterogeneous integration of smaller chiplets using advanced packaging has emerged as a promising paradigm for addressing this limit and meeting performance demands. These approaches offer a significant cost reduction and higher manufacturing yield than monolithic 2D integrated circuits. However, the compact arrangement and high compute density of these systems exacerbate thermal management challenges, potentially compromising performance. Addressing these thermal modeling challenges is critical, especially as system sizes grow and different design stages require varying levels of accuracy and speed. Since no single thermal modeling technique meets all these needs, this paper introduces MFIT, a range of multi-fidelity thermal models that effectively balance accuracy and speed. These multi-fidelity models can enable efficient design space exploration and runtime thermal management. Our extensive testing on systems with 16, 36, and 64 2.5D integrated chiplets and 16$\times$3 3D integrated chiplets demonstrates that these models can reduce execution times from days to mere seconds and milliseconds with negligible loss in accuracy.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXX
1. Introduction
@@
Massive data from different modalities, including text, images, video, and speech, are continuously produced by various sensors. At the same time, increasingly complex artificial intelligence (AI) and machine learning (ML) algorithms process this data to enable new applications that were previously impractical. This trend dictates the design of large-scale chips with high memory and compute capabilities, offering a high degree of parallelism (src, 2021; Association, [n. d.]). Traditional 2D chip design and packaging technologies cannot sustain this need due to the low yield of large monolithic planar chips and the corresponding increase in fabrication cost (Cheng et al., 2020). Therefore, new design approaches are required to meet the increasing demand for computing power and memory capacity (src, 2021).
@@
2.5D and 3D chiplet-based architectures have emerged as promising alternatives to traditional monolithic 2D chips due to their lower fabrication costs (Shao et al., 2021; Sharma et al., 2023; Krishnan et al., 2022). Compared to conventional monolithic systems, chiplet-based systems integrate multiple small pre-fabricated chips (chiplets) on a silicon interposer, which facilitates data exchange, as illustrated in Fig. 1(a). 3D packaged systems expand on this approach by stacking multiple chiplets vertically and connecting them with vertical vias, creating a more compact system as illustrated in Fig. 1(b). The smaller size of these chiplets enables a higher yield and lower overall manufacturing cost than traditional monolithic dies (Stow et al., 2016). Additionally, this modular approach facilitates scaling the system sizes and enables heterogeneous integration of different chiplet types, e.g., memory, processing, and processing-in-memory chiplets. Hence, emerging 2.5D and 3D architectures enable a new cost-effective avenue for compact scale-out implementations of various emerging compute- and data-intensive applications, including AI/ML. Indeed, these advantages have led to industrial adoption by companies including Intel (Won, 2023; Black et al., 2006), AMD (Naffziger et al., 2021; Bharadwaj et al., 2020; Agarwal et al., 2022), and NVIDIA (Shao et al., 2021).
@@
(a) A 16 - 2.5D integrated chiplet-based system. The magnified view on the right shows a detailed structure of a single chiplet.
@@
(b) A 16$\times$3 - 3D integrated chiplet-based system. The magnified view on the right shows a detailed structure of three chiplets stacked.
@@
Figure 1. 2.5D/3D integrated chiplet systems considered in this work, showing the chiplets, interposer, and part of the substrate.
Thermal bottlenecks have long been a significant barrier to increasing the performance of computing systems. 2.5D and 3D integrated systems exacerbate this barrier due to their dense integration and unique physical structure (Park et al., 2024). Unlike monolithic chips, where heat spreads uniformly across a continuous die, 2.5D chiplet-based systems conduct heat laterally through the interposer and vertically through the heat spreader. Although the interposer, often built with materials of relatively high thermal conductivity, can help alleviate localized thermal hotspots, chiplet-based designs introduce new sources of thermal stress. In particular, a single package can accommodate more active silicon by integrating multiple smaller chiplets, thereby increasing the total compute throughput and system power. If the total power consumption increases faster than the area, the package-level power density also becomes larger than a monolithic chip. Furthermore, while the interposer aids in thermal spreading, it also hosts numerous communication links whose resistive Joule heating contributes to additional thermal challenge (Ma et al., 2024). Likewise, heat also flows vertically between adjacent stacked chiplets in a 3D chiplet-based system.
@@
Figure 2. Summary of the multi-fidelity thermal models. (1) Fine-grained FEM models capture precise geometry but are too complex to simulate the entire chiplet-based system. (2) Abstracted FEM models are derived from the fine-grained model to simulate large-scale systems with negligible impact on accuracy ($<\mathrm{0.5}^{\circ}$, which is 0.5–1% around the temperatures of interest). (3) Since abstract FEM models are still too slow for DSE, they are used to tune thermal RC circuit models, which introduce less than 1.7C (1–3.5% around the temperatures of interest) error). (4) Further abstraction reduces the execution time to milliseconds using DSS models developed for specific system configurations, enabling runtime thermal management.
@@
The semiconductor chip design cycle spans multiple phases: system specification, architecture exploration, logic design, physical design and validation, fabrication, and post-silicon optimization/validation. Each phase has a unique set of design constraints and requirements. For example, the lack of a test chip during the pre-silicon phases requires simulation and analytical models. Finite Element Method (FEM) simulations offer the most accurate approach for pre-silicon thermal analysis (Sultan et al., 2019). They can serve as a reference and enable heat flow studies to guide the design process. However, they are too slow for practical architecture and design space exploration (DSE), as illustrated in Fig. 2. Modeling the package as a thermal RC (resistive-capacitive) network can significantly accelerate simulations with acceptable accuracy loss (Skadron and otehrs, 2004; Sridhar et al., 2013). Since each node in the thermal circuit corresponds to a specific location in the package, thermal RC models solve discretized versions of the FEM models in space. Hence, they enable thermally-aware DSE and optimization with a finite number of discrete hotspot nodes. However, the thermal resistance/capacitance values and the circuit topology must accurately reflect the chip geometry and material properties for reliable results. Since the thermal RC models solve continuous-time ordinary differential equations (ODEs), they have execution times in the order of seconds to minutes. Therefore, they cannot be used for runtime optimization tasks such as dynamic thermal and power management (DTPM). With a given sampling period, one can discretize them in the time domain. The resulting discrete state-space (DSS) models significantly reduce runtime at the cost of further abstracting the model from the physical package. Consequently, they apply only to the specific configurations for which they are developed.
@@
There is a strong need for tools to accurately analyze the thermal behavior of 2.5D and 3D integrated systems and guide their design process. However, no single modeling technique can alone address the needs of all design phases. To address this critical gap, this paper proposes MFIT, a multi-fidelity thermal modeling framework that synergistically exploits the strengths of each class of models (FEM, thermal RC, and DSS). We use this framework to produce a set of thermal models that can guide the entire design cycle, unlike a point solution that can serve a specific portion of the design process. The elements of this set not only cover complementary parts of the design cycle but support each other and produce consistent results. We first develop a fine-grained FEM model of the target package as a reference. Since it is slow and computationally expensive, we next judiciously design an abstracted version of this fine-grained FEM model to simulate an entire package in days while maintaining accuracy. To enable fast DSE, MFIT also incorporates thermal RC circuit models verified against the reference FEM models. Our thermal RC models run in the order of seconds while leading to less than 1.7C error when the package temperature is around 100C, as summarized in Fig. 2. Hence, they can be used for pre-silicon architectural optimization, such as mapping the workloads to chiplets, network-on-interposer design, and chiplet placement for 2.5D and 3D stacked systems. Finally, MFIT derives a final class of models by discretizing the thermal RC models, enabling runtime thermal management and large-scale DSE in the order of milliseconds. Since these discrete models are generated for a specific sampling period and configuration, they are regenerated from the RC model if the target configuration changes automatically in a few milliseconds. In summary, we obtain a set of multi-fidelity thermal models that guide and complement each other to cover all design phases.
@@
  • A family of open-source multi-fidelity thermal models that span a broad accuracy (reference to 1.7C) and speed (days to milliseconds) range,
@@
  • Open-sourced code for thermal RC and DSS models at github.com/AlishKanani/MFIT. The FEM models created for this work are included as well.
@@
The remainder of the paper is organized as follows. Section 2 and Section 3 discuss related work and background on FEM. Section 4 presents the proposed multi-fidelity thermal modeling framework. Finally, Section 5 presents the experimental evaluation, and Section 6 concludes the paper.
2. Related Work
@@
2.5D and 3D integration-based systems are becoming mainstream due to higher performance and lower manufacturing costs than monolithic chips. Both domain-specific and general-purpose 2.5D and 3D architectures have been explored to date (Bharadwaj et al., 2020; Krishnan et al., 2021; Sharma et al., 2022a, 2023; Shao et al., 2021; Eckert et al., 2014; Agarwal et al., 2022). SIMBA is one of the first prototype multi-chip modules with 36 chiplets designed for inference with deep models (Shao et al., 2021). Similarly, Floret is a data center-scale architecture for accelerating convolutional neural network (CNN) inference tasks by exploiting dataflow patterns (Sharma et al., 2023). Loi et al. (Loi et al., 2006) analyze the performance benefits of vertically integrated (3D) processor-memory hierarchy under thermal constraints. Similarly, Eckert et al. (Eckert et al., 2014) consider processing-in-memory (PIM) architectures implemented using 3D die stacking. They study the thermal constraints across different processor organizations and cooling solutions to identify viable solutions. Our proposed open-source thermal models catalyze similar thermal analysis and optimization studies for 2.5D and 3D integrated systems.
@@
The most accurate and direct thermal evaluation approach is temperature measurements on a hardware system. It can be performed using thermal imaging (Sadiqbatcha et al., 2019; Amrouch and Henkel, 2015) or temperature sensors (Zhang et al., 2021, 2008). However, the availability of the target hardware is a significant limitation. For example, large-scale 2.5D and 3D chiplet systems with tens of chiplets do not exist yet, while smaller prototypes and commercial systems provide limited insights applicable for larger systems (Naffziger et al., 2021; Zaruba et al., 2020). This limitation motivated FEM-based modeling as the most accurate way to analyze the heat flow and temperature (Sultan et al., 2019). Proprietary software, such as ANSYS Fluent (Matsson, 2023) and COMSOL (COMSOL, [n. d.]), are commonly used for FEM simulations. Since FEM suffers from computational cost, detailed FEM solutions are suitable only for small designs and validating analytical models (Matsson, 2023). For example, the authors of (Zhou et al., 2022) employ FEM to simulate a two-chiplet system on an interposer. They employ abstracted FEM models for both $\mu$-bumps and C4 bumps to speed up the process, effectively reducing computational complexity before tackling the entire system.
@@
The computational overhead and impractically high execution time of FEM solvers motivate analytical models that enable rapid thermal evaluation in early design phases. The most common method involves constructing thermal RC networks and solving the corresponding system of ODEs. Popular thermal simulators such as HotSpot (Skadron and otehrs, 2004) leverage this method, focusing on the microarchitectural layout blocks to facilitate design space exploration and early-stage thermal-aware layout and placement. Similarly, 3D-ICE (Sridhar et al., 2013) models liquid cooling with microchannels embedded between silicon layers. PACT (Yuan et al., 2021) also employs a similar methodology by utilizing SPICE tools as solvers, focusing on standard-cell-level thermal analysis for 2.5D systems. However, these tools lack essential features required for accurately modeling chiplet-based heterogeneous systems and are not fast enough for large-scale, thermally-aware design space exploration (DSE) of multi-chiplet 2.5D and 3D integrated systems, as summarized in Table 1 and detailed next.
Table 1. Qualitative analysis of existing analytical thermal models and our proposed thermal RC and DSS models. A detailed quantitative runtime comparison is completed in Section 5.
Non-uniform
grid
Anisotropic
materials
Non-homogeneous
layers
Heat dissipation from both
boundaries of the package
Flexible with
Architecture changes
Hotspot (Skadron et al., 2004) $\times$ $\times$
PACT (Yuan et al., 2021) $\times$ $\times$ $\times$
3D-ICE (Sridhar et al., 2013) $\times$ $\times$ $\times$ [^1] With Non-uniform grid, 3D-ICE does not support secondary heat flow path.
Thermal RC
DSS $\times$
2.5D and 3D packages often involve materials with varying thermal conductivity across different directions. For example, the thermal conductivity of the C4 layer is higher in the vertical direction than in the lateral direction. The existing thermal models (Skadron and otehrs, 2004; Sridhar et al., 2013; Yuan et al., 2021) do not account for these anisotropic properties. Moreover,  (Skadron and otehrs, 2004; Yuan et al., 2021) assume a uniform grid size for all material layers (e.g., interposer, C4 bumps, chiplets). Non-uniform grids are useful for simulating regions with substantial spatial variation in temperature without unnecessarily increasing computational costs for simulating regions of moderate temperature variation. While 3D-ICE supports non-uniform grids for 3D architectures, it does not allow non-homogeneous layers, where different regions within the same layer can have distinct material properties. This limitation is particularly significant for emerging technologies, such as active interposer-based systems where chiplets are embedded (Vanna-Iampikul et al., 2023). In contrast, our multi-fidelity models address these limitations by enabling variable thermal conductivity across different directions and allowing flexible grid sizes for each layer and block. Additionally, our models support heat dissipation from both package boundaries, a crucial aspect overlooked by some existing approaches. This capability is essential for accurately modeling advanced cooling techniques, such as immersion cooling systems and flip-chip packages, which rely on efficient heat dissipation through both package boundaries to maintain optimal performance and reliability (Li et al., 2010; Huang et al., 2023). Table 1 provides a detailed qualitative comparison of existing thermal modeling tools. Detailed quantitative runtime comparisons between MFIT and state-of-the-art approaches are presented in Section 5.
Architecture-level thermal RC models are commonly used for offline studies. For example, temperature sensor placement requires using a thermal model (Zanini et al., 2010; Chundi et al., 2017). Similarly, thermally-aware chiplet placement techniques in (Ma et al., 2021; Eris et al., 2018; Li et al., 2022) rely on the use of architectural thermal simulators to optimize the placement of chiplets and identify possible hotspots. While existing simulators can provide the functionality required for these studies, their prohibitive runtime prevents more extensive evaluations from being performed. Faster thermal simulation tools that do not sacrifice accuracy are required to enable more comprehensive work.
@@
However, existing architectural simulators are inadequate for dynamic thermal and power management (DTPM). DTPM requires a much faster temperature estimation time, on the order of milliseconds, for real-time temperature management (Sharifi and Rosing, 2010; Bhat et al., 2018). DSS models address this need by deriving a discrete-time linear time-invariant system that models the thermal dynamics at fixed locations as a function of the power consumption. For instance, TILTS (Han et al., 2007) discretizes the power inputs to the chip over fixed time intervals to accelerate thermal simulations. Hence, it needs to be reproduced when the timing requirements or the underlying hardware configuration change. The speedup gain offsets the loss of the explicit connection to the hardware parameters (e.g., thermal conductance and capacitance) and generality.
@@
The results of FEM, thermal RC, or other thermal simulations can also be used to train physics-informed machine learning techniques to model heat transfer in integrated circuits to reduce the thermal modeling effort (Cai et al., 2021). For example, a recent technique collects data from numeric simulations and trains a random forest model to predict the convection heat transfer coefficients for a nonlinear heat transfer problem (Kwon et al., 2020). Similarly, Hwang et al., (Hwang et al., 2018) present closed-form models derived from numerical simulations for tapered micro-channels to analyze the heat transfer performance as a function of the channel geometry. In contrast to individual classes of thermal models, this work proposes a framework to produce a family of multi-fidelity thermal models for 2.5D and 3D chiplet-based systems. The specific set of models designed with this framework covers a wide range of accuracy and execution time trade-offs, making them suitable for different design phases. Additionally, they can be augmented by additional models, such as physics-informed ML models, with complementary accuracy and execution time trade-offs.
3. Finite Element Method for Thermal Analysis